Mrs. Simpson buys loaves of bread and quarts of milk each week at prices of $3 and $2.50, respectively. At present she is buying these two products in amounts such that the marginal utilities from the last units purchased of the two products are 80 and 70 utils, respectively. Is she buying the utility maximizing combination of bread and milk? If not, how should she reallocate her expenditures between the two goods?
1
Expert's answer
2020-10-01T09:59:04-0400

The utility maximizing combination of bread and milk is consumed when:

"MUb\/Pb = MUm\/Pm,"

At present MUb/Pb = 80/3 = 26.67,

MUm/Pm = 70/2.5 = 28.

So, the consumption is not optimal, and he should consume less bread and more milk.